In 1862, two companies were granted the rights to build the transcontinental railroad from Omaha, Nebraska, to Sacremento, California. The Central Pacific railroad began in 1863 from Sacremento heading east. The Union Pacific began 24 months later, leaving from Omaha and heading west. The Central Pacific averaged 8.75 miles of track each month, while the Union Pacific averaged 20 miles of track per month. The two companies met at Promontory, Utah, after completing 1590 miles of track. How much track did each build?
Answer by ankor@dixie-net.com(22740)

In 1862, two companies were granted the rights to build the transcontinental railroad from Omaha, Nebraska, to Sacramento, California.
The Central Pacific railroad began in 1863 from Sacramento heading east.
The Union Pacific began 24 months later, leaving from Omaha and heading west.
The Central Pacific averaged 8.75 miles of track each month, while the Union Pacific averaged 20 miles of track per month.
The two companies met at Promontory, Utah, after completing 1590 miles of track. How much track did each build?
:
First find out how many months to complete the task
:
m = no. of months of work by the Union Pacific
then
(m+24) = no. of months of work of the Central Pacific
:
20m + 8.75(m+24) = 1590
20m + 8.75m + 210 = 1590
28.75m = 1590 - 210
28.75m = 1380
m = 1380/28.75
m = 48 month of work by the UP
then
48(20) 960 mi of track completed by the UP
:
and
8.75(48+24) = 630 mi of track completed by the CP
;
:
See if that adds up: 960 + 630 = 1590 total miles